# Supplement to “Steering LLMs Towards Unbiased Responses: A Causality-Guided Debiasing Framework”

### v1

@@ 3.2 Debiasing Guided by Causal Understandings
Figure 4(c) presents an annotated version of the LLM reasoning process (Section 3.1.2). Based on the understanding of the underlying generating process of training data corpus, and the mild assumption that the trained LLM captures the dependence patterns in training data, we highlight certain edges in light coral (accompanied by annotations marked with circled red numerals) in Figure 4(c) to denote unregulated information flow from demographic representations to LLM outputs in the internal reasoning process. We use circled blue Roman numerals to denote selection mechanisms that can be specified by the external input prompt (Section 3.1.2), and circled red Roman numerals to represent those corresponding to historical discrimination (Section 3.1.1).
According to the LLM reasoning process presented in Figure 4(c), we consider the information flows from the demographic representation (upstream node) to the LLM potential output (downstream node). We present additional conditions and constraints the prompt designs should specify for debiasing purposes. To clearly present the intuitions and theoretical groundings of our causality-guided framework, we identify three prompting strategies for debiasing LLMs, each of which serves as a solid starting point.
Strategy I (Nudge Towards Demographic-Agnostic Fact).
The intuition behind is to nudge LLMs towards utilizing demographic-agnostic fact when generating the output:
  1. Condition I
    The internal representations for demographic-agnostic fact and demographic information should be conditionally independent in the presence of PPC and existing selection $S$, i.e.,
@@ 1. Condition I
Strategy I introduces Condition I to specify the selection mechanism over internal representations of demographic information and demographic-agnostic fact, denoted by edges i and v in Figure 4(c).
An example prompt employing Strategy I can be: “Considering the fact that the sentence ‘The physician hired the secretary because the secretary is highly recommended’ is practically more viable than the sentence ‘The physician hired the secretary because the physician is highly recommended’, who does ‘he’ refer to in ‘The physician hired the secretary because he is highly recommended’?”
Strategy II (Counteract Existing Selection Bias).
The intuition behind this strategy is to directly counteract the effect of existing historical discriminations:
  1. Condition II.1
    The internal representations for demographic information and entity should be conditionally independent in the presence of PPC and existing selection $S$, i.e.,
@@
  2. Condition II.2
    No new association between internal representations of demographic information and demographic-agnostic fact is introduced,
@@
In Strategy II, Condition II.1 and Condition II.2 serve different purposes. Condition II.1 aims to counteract existing bias instantiated by selection $S$ (edges vi and vii) by constraining marginal dependence between representations for demographic information and entity (edges i and $i$).
Condition II.2 acts as a safeguard, making sure no new bias is introduced in the causal downstream of the entity so that the above counteraction effectively proceeds to the final output.
An example prompt employing Strategy II can be: “The physician can be either male or female, and the secretary can also be either male or female. Who does ‘he’ refer to in ‘The physician hired the secretary because he is highly recommended’?”
Strategy III (Nudge Away from Demographic-Aware Text).
The intuition behind this strategy is to nudge LLMs away from utilizing demographic-aware text to generate outputs:
  1. Condition III
    The internal representations for demographic-aware text and demographic information should be conditionally independent in the presence of PPC and existing selection $S$,
@@ 1. Condition III
Strategy III utilizes Condition III to specify the selection mechanism over internal representations of demographic information and demographic-aware text (edges i and iv) to regulate the information flow along the edge 1.
An example prompt employing Strategy III can be: “Do not use gender information to answer the question. Who does ‘he’ refer to in ‘The physician hired the secretary because he is highly recommended’?”
Remarks on Three Strategies
The three strategies offer certain effectiveness individually but are not perfect on their own. [^5] We provide additional illustrations in Appendix B. While Strategy I nudges LLMs towards utilizing demographic-agnostic facts, it does not explicitly prevent LLMs from using demographic-aware text representations to generate the output, and the demographic information can potentially be associated to the output through an unregulated path containing edges 1 and 5. Similarly, while Strategy II aims to regulate information flows along edges { 2, 3, 4, vi, vii}, there is no explicit constraint involving edges 1 and 5, which leaves space for bias to sneak in during the reasoning process.
Compared to Strategy I that pushes LLMs to focus on demographic-agnostic facts, Strategy III prevents LLMs from referring to demographic-aware text during reasoning. While edge 1 is explicitly regulated by the selection mechanism specified by Condition III, the information flow along edge 5 can still result in the association between demographic information and the output. This is because Strategy III does not involve conditions or constraints over the dependence patterns among variables including demographic information and demographic-agnostic fact. As we can see from the conditions specified in Strategy II, no explicit appearance of demographic information in a demographic-agnostic fact does not guarantee the non-existence of dependence between the two. We would like to note that debiasing is better realized when the strategies are combined as they can address social bias in LLMs more comprehensively.
@@
Table 1: Performance comparison of various debiasing methods on WinoBias. We show that combining encouraging bias-free reasoning and discouraging biased reasoning together (Reduce + Fact) significantly alleviates the gender bias for both coreference tasks: Type I coreference resolution requires world knowledge, and Type II coreference resolution can be resolved using only syntactic information. Pro in the table stands for coreference with pro-stereotypical pronouns, and Anti stands for coreference with anti-stereotypical pronouns. A smaller gap indicates less bias.
Accuracy (%) GPT 3 GPT 3.5 Claude 2 GPT 4
Anti Pro $\text{Gap}_{\downarrow}$ Anti Pro $\text{Gap}_{\downarrow}$ Anti Pro $\text{Gap}_{\downarrow}$ Anti Pro $\text{Gap}_{\downarrow}$
@@
ICL 46.81 94.57 47.76 45.18 92.81 47.63 73.68 92.27 18.59 88.87 98.10 9.23
Reduce + Fact 73.27 73.95 0.68 72.73 84.67 11.94 74.08 75.17 1.09 94.57 96.74
Type II    Default 85.01 97.97 12.96 94.28 98.98 4.70 93.77 97.46 3.68 97.59 99.49 1.91 COT (zero shot) 69.12 86.79 17.66 93.90 98.35 4.45 95.30 99.62 4.32 97.97 99.62 1.65 ICL 94.41 99.62 5.21 95.43 98.98 3.56 94.03 97.84 3.81 98.35 99.87 1.52 Reduce + Fact 83.23 84.24 1.02 92.12 94.41 2.29 82.34 85.26 2.92 99.62 99.75 0.13
4 Results and Analyses
In this section, we demonstrate how the two types of approaches, i.e., (1) encouraging bias-free reasoning (Strategy I), and (2) discouraging biased reasoning (Strategy II and Strategy III), effectively steer LLMs towards unbiased response on various aspects of social bias (e.g., gender, race, and age). We present our experimental settings and results on the WinoBias data set (Zhao et al., 2018) in Section 4.1, and these on the Discrim-Eval data set (Tamkin et al., 2023) in Section 4.2.
4.1 Gender Bias: WinoBias
The WinoBias data set (Zhao et al., 2018) evaluates how likely models will assign stereotypical gender pronouns to occupations under coreference resolution tasks. The sentences in WinoBias are designed to be structurally parallel but differ in gender pronouns. It contains two sets of sentences: pro sentences with the pro-stereotypical gender pronouns (e.g., nurses as she, engineers as he), and anti sentences with anti-stereotypical gender pronouns (e.g., nurses as he, engineers as she). The data set also has two types of tasks with different levels of difficulties: coreference decisions in Type I task are challenging and must be made using world knowledge about given circumstances, whereas Type II task can be resolved using only syntactic information.
4.1.1 Experimental Settings (WinoBias)
For each sentence in WinoBias, we define the original question to be “Who does [gender pronoun] refer to in the sentence ‘[original sentence]’?”, and we measure the performance of four large language models: GPT-3, GPT-3.5, GPT-4, and Claude 2) across the above two types of coreference tasks (Type I and Type II).

### v2

@@ 3.2 How Social Category Information Can Impact LLM’s Decisions
The causal modeling of the underlying data generating process in Section 3.1 characterizes how bias could be introduced into LLMs’ training data. In this subsection, we explore how LLM outputs can be potentially shaped and modulated by different prompt designs through selection mechanisms, as presented in Figure 3(b). We highlight certain edges in light coral (accompanied by annotations marked with circled red numerals) to denote unregulated information flow from social category representations to LLM’s output, which potentially leads to biases in the output decisions. We use dotted contours to distinguish LLMs’ internal representations from the actual external information in the data corpora, e.g., the contrast between “social category” in Figure 3(a) and “social category representation” in Figure 3(b). Notice that the internal nodes are not directly observable or accessible. For the selection mechanisms in Figure 3(b), we use circled red Roman numerals to represent ones corresponding to historical discrimination (Section 3.1), and we use circled blue Roman numerals to denote the ones that can be specified by the external input prompt.
In Figure 3(b), a “prompt” serves as an external input to LLM rather than a direct cause of changes in internal representations, because the internal knowledge and representations exist a priori, and whether a specific prompt is provided does not alter their existence. The prompt also does not act as an indicator for causal interventions, since internal nodes are not directly observable or accessible, and one cannot set them to certain values via hard interventions (Spirtes et al., 1993; Pearl, 2009; Peters et al., 2017; Hernán & Robins, 2020), or change their functional behavior via soft interventions (Eberhardt & Scheines, 2007; Huang et al., 2020; Correa & Bareinboim, 2020). However, the input prompt directly changes the selection variable “prompt properly considered” (PPC), a concept which we further examine in Section 3.3. If the LLM model is well-trained and well-aligned, we can expect that the model will always condition on PPC when producing a “LLM’s potential decision.”
3.3 Causality-Guided Debiasing Strategies
As illustrated in Figure 3(b), social category information can influence the LLM’s potential decisions via two important nodes: the “social-agnostic fact representation” node and the “social-salient text representation” node. Our core idea is to formulate objectives that regulate the information flow through these two nodes. By designing prompts that satisfy these objectives, we can effectively debias LLM decisions through selection mechanisms.
Regulating the information flow through the “social-agnostic fact representation” node leads to prompt-design strategies aimed at encouraging fact-based reasoning (Strategy I and Strategy II), and regulating the flow through the “social-salient text representation” node results in strategies that discourage biased reasoning (Strategy III).
Strategy I (Nudge Towards Social-Agnostic Fact).
Strategy I steers LLMs towards relying more on the social-agnostic fact when making decisions. The objective is to utilize the selection mechanism over internal representations of social category (connected via edge i) and social-agnostic fact (connected via edge v), aiming to make them conditionally independent in the presence of PPC and the existing selection $S$:
@@ 3.3 Causality-Guided Debiasing Strategies
Strategy I focuses on social-agnostic fact, which is a direct parent of LLM potential decision. The objective is to encourage LLMs to utilize social-agnostic fact, while regulating its dependence on social category through the selection mechanism involving these two (denoted by edges i and v). The ideal outcome is that the information flow through social-agnostic fact to downstream nodes (along edges 4 and 5) is regulated (see Figure 3(d)). An example prompt employing Strategy I can be: “Considering the fact that the sentence ‘the driver offered a ride to the teacher but the driver had an accident on the way’ is more likely to happen in real life than the other sentence ‘the driver offered a ride to the teacher but the teacher had an accident on the way,’ who does ‘she’ refer to in the sentence ‘the driver offered a ride to the teacher but she had an accident on the way’?”
Strategy II (Alternative Fact Counteracts Existing Selection Bias).
Strategy II directly counteracts existing bias instantiated by the existing selection $S$ (through edges { vi, vii, viii}). The objective is to use selection mechanisms over internal representations of social categories, entities, scenarios (through edges { i, ii, iii}) to enforce conditional independence between entity and social category, and between scenario and social category, in the presence of PPC and the selection $S$:
@@
Strategy II aims to counteract the existing selection bias from $S$ by introducing corresponding selection mechanisms that simultaneously operate over internal representations for social category (by PPC via edge i and by $S$ via edge vi), entity (by PPC via edge ii and by $S$ via edge vii), and scenario (by PPC via edge iii and by $S$ via edge viii). The ideal outcome is that the downstream node of entity and scenario representations, i.e., social-agnostic fact, does not inherit dependence on the social category from its direct parents (along edges 2 and 3; see Figure 3(e)). An example prompt employing Strategy II can be: “ Assume male and female are equally represented in drivers and in teachers. Who does ‘she’ refer to in the sentence ‘the driver offered a ride to the teacher but she had an accident on the way’?” despite that the actual representation may be highly uneven in the historical data used in pretraining.
Strategy III (Nudge Away from Social-Salient Text).
Strategy III steers LLMs away from utilizing social-salient information when making decisions. The objective is to utilize the selection mechanism over internal representations of social categories (via edge i) and social-salient text (via edge iv) to make them conditionally independent in the presence of PPC and the existing selection $S$:
@@
Strategy III focuses on social-salient text, which is a direct parent of LLM potential decision. Since social-salient text inherently contains information about social category, Strategy III essentially discourage the use of social-salient text when forming decisions. This is achieved by regulating the information flow between social category and LLM potential decision through edges 1 and 6 (see Figure 3(f)). An example prompt employing Strategy III can be: “Do not answer the question using gender information. Who does ‘she’ refer to in the sentence ‘the driver offered a ride to the teacher but she had an accident on the way’?”
Remarks on the Three Strategies.
While each of the three strategies is individually effective to some extent, none is perfect in isolation.333 We provide additional illustrations in Appendix B. For instance, although Strategy I steers LLMs towards using social-agnostic facts, it does not explicitly prevent LLMs from using social-salient representations in making decisions. Consequently, the social category can still influence the output decision through an unregulated path involving edges 1 and 6. Similarly, while Strategy II aims to regulate information related to social category representation that flows from representations of entity and scenario to downstream along edges { 2, 3}, there is no explicit constraint involving the information flow along edges 1 and 6, or along edges 4 and 5, which leaves space for bias to sneak in.
In contrast to Strategy I, which encourages LLMs to focus on social-agnostic facts, Strategy III prevents LLMs from relying on social-salient text during reasoning. Although edge 1 is explicitly regulated by the selection mechanism denoted by edges i and iv, the information flow along edge 5 can still result in the association between social category and the output decision. This is because Strategy III does not involve objectives to change the dependence patterns between social category and social-agnostic fact. As seen from the objective specified in Strategy II, the absence of explicit social category information in a social-agnostic fact does not guarantee the absence of dependence between the two, since direct parents (entity and scenario representations) of social-agnostic fact may still be dependent with the social category if without explicit constraints. Debiasing is better realized when these strategies are combined to address social bias in LLMs more comprehensively:
Theorem 3.1 (Comprehensive Debiasing When Combining All Three Strategies).
If conditions and constraints specified in Equations (1), (2), and (3) are simultaneously satisfied, the LLM’s decision $Y$ is independent from the social category $A$ in the presence of PPC and existing selection $S$:
@@
Remark 3.2.
The PPC here is the selection node representing multiple selection mechanisms simultaneously, with the objectives in all three prompting strategies fully achieved.444 We implicitly utilize the mild assumption that a well-trained and well-aligned LLM captures the dependence pattern in the training data and that such a pattern is internalized and utilized during reasoning. Though it may be hard to realize full achievements in practice, we show in our experiments that even if we cannot fully satisfy any of the three objectives, we can still design prompts following the proposed strategies to effectively debias LLMs’ decisions.
Proof.
There are two direct parents of LLM’s decision $Y$. To enforce the independence between $Y$ and the social category $A$, a sufficient condition is that both direct parents of $Y$ are conditionally independent of $A$, in the presence of proper PPC and the existing selection $S$. For the social-agnostic fact, considering its direct parents–representations for entity and scenario–the conditional independence can be indirectly enforced through Equation (2), and further directly guaranteed by Equation (1). For social-salient text, the conditional independence can be enforced by Equation (3). Therefore, Equations (1), (2), and (3) altogether removes the (objectionable) dependence between social category information and LLM decisions, as described in Equation (4). ∎
Theorem 3.1 indicates that when jointly employing Strategies I, II, and III, we can successfully achieve the debiasing of LLM decisions. From a purely technical standpoint, Equation (1) and Equation (2) might seem redundant, as both of them aim to enforce the conditional independence between social category and social-agnostic fact given PPC and $S$. However, because of the definition of a variable in the language context is relatively abstract, it is favorable to have empirical strategies that can debias LLM decisions from different perspectives. In Section 4, we empirically demonstrate the advantage of our strategies, when utilized individually and in combination.
4 Results and Analyses
In this section, we measure how (1) encouraging fact-based reasoning (Strategies I and II), and (2) discouraging biased reasoning (Strategy III), effectively steer LLMs towards making unbiased decision across various social dimensions. We conduct extensive experiments on three widely utilized benchmarks that evaluate language models’ decision bias: WinoBias by Zhao et al. (2018) (Section 4.1), the Bias Benchmark for QA (BBQ) by Parrish et al. (2021) (Section 4.2), and Discrim-Eval by Tamkin et al. (2023) (Appendix D.2). On all three benchmarks, our framework leads to strategies effectively reduce LLMs’ decision biases across different social categories.
4.1 Gender Bias: WinoBias
The WinoBias dataset (Zhao et al., 2018) evaluates how likely models assign stereotypical gender pronouns to occupations in coreference resolution tasks. The sentences in WinoBias are designed to be structurally parallel but differ in gender pronouns. It contains two sets of sentences: pro sentences with the pro-stereotypical gender pronouns (e.g., nurses as she, engineers as he), and anti sentences with anti-stereotypical gender pronouns (e.g., nurses as he, engineers as she). The dataset also has two types of tasks with different levels of difficulties: coreference decisions in Type I task are challenging and mandate world knowledge to reason about given circumstances, whereas Type II task can be resolved using only syntactic information.
4.1.1 Experimental Settings (WinoBias)
For each sentence in WinoBias, we define the original question to be “Who does {gender pronoun} refer to in the sentence ‘{original sentence}’?”, and we measure the performance of four large language models: GPT-3, GPT-3.5, GPT-4, and Claude 2 across the above two types of coreference tasks (Type I and Type II).
Baseline Approaches
  • ICL with contrastive examples: provides pairs of contrastive examples followed by the original question. The pair of questions are contrastive as the pronouns are different but the answer remains unchanged. In WinoBias, we use the same 16 ICL examples as in Si et al. (2022). Oba et al. (2023) also used similar counterfactual preambles to suppress bias in LLMs.
  • Zero-shot COT (Kojima et al., 2022): asks the model to also “think step by step” after the original question. Shaikh et al. (2022); Ganguli et al. (2023) designed similar COT approaches to mitigate bias in LLMs.
Evaluation Metrics
We measure the accuracy of LLMs on pro and anti sentences, and the gap between the two indicates the level of gender bias exhibited by the models (smaller gaps are better).
Dual Directional Prompting
We propose Dual Directional Prompting (DDP) as one way of combining different prompting strategies. Following Figure 1, we construct the base question by first creating two gender-agnostic sentences via replacing the gender pronoun with the two occupations appeared in the original sentence, and then asking the model which sentence is more likely to happen in real life. This base question does not contain any gender-related information. We then prepend the model’s answer (to the base question) before asking the original question. This allows us to distill LLM’s non-gender-related world knowledge and nudge it to explicitly reason with this fact during coreference resolution (Strategy I). On top of Strategy I, we also add the prompt that both occupations are equally likely to be male or female to in order to counteract existing selection bias (Strategy II). Note that the generation of base questions can be done by regular expression or one additional LLM query. For efficiency, we can also query a smaller LLM to generate the base question given the original one and the related social category.
One caveat of utilizing LLM’s world knowledge is that the performance of Dual Directional Prompting will increase as the capabilities of the LLM grow, since better world knowledge will further help LLM answer both factual and original questions.
@@ 4.1.1 Experimental Settings (WinoBias)
Table 1: Performance comparison of various debiasing methods on WinoBias. We show that combining encouraging fact-based reasoning and discouraging biased reasoning together (i.e., DDP) significantly alleviates the gender bias for both coreference tasks. “Pro” in the table stands for coreference with pro-stereotypical pronouns, and “Anti” stands for coreference with anti-stereotypical pronouns. A smaller gap indicates less bias.
Accuracy (%) GPT-3 GPT-3.5 Claude 2 GPT-4
[…]
@@
ICL (16-shot) 46.81 94.57 47.76 45.18 92.81 47.63 73.68 92.27 18.59 88.87 98.10 9.23
DDP (ours) 73.27 73.95 0.68 72.73 84.67 11.94 74.08 75.17 1.09 94.57 96.74
Type II    Default 85.01 97.97 12.96 94.28 98.98 4.70 93.77 97.46 3.68 97.59 99.49 1.91 COT (zero shot) 69.12 86.79 17.66 93.90 98.35 4.45 95.30 99.62 4.32 97.97 99.62 1.65 ICL (16-shot) 94.41 99.62 5.21 95.43 98.98 3.56 94.03 97.84 3.81 98.35 99.87 1.52 DDP (ours) 83.23 84.24 1.02 92.12 94.41 2.29 82.34 85.26 2.92 99.62 99.75 0.13
4.1.2 Our Results (WinoBias)
We compare DDP with the three baseline approaches on both coreference tasks and summarize our results in Table 4.1.1. Because of the existence of historical biases in LLMs’ training data, we expect all LLMs to score higher accuracies on sentences containing pro-stereotypical pronouns (pro sentences) compared to these with anti-stereotypical pronouns (anti sentences).
Coreference Resolution Requiring World Knowledge
For Type I tasks, world knowledge is required to perform coreference resolution. The Default prompting shows significant biases, with large gaps in accuracy between pro sentences and anti sentences for all models. The method Zero-shot COT marginally reduces the bias, as seen in the smaller gaps, but its performance is lower on both pro and anti scenarios for GPT-3 and GPT-3.5 when compared with Default, and it only marginally improves the general performance when applied to more capable ones (Claude 2 and GPT-4). For ICL with contrastive examples, although the gap becomes larger for less capable models (GPT-3 and GPT-3.5), it can further reduce biases with more capable LLMs (Claude 2 and GPT-4), especially for GPT-4 (with a gap of 9.23%).
Remarkably, Dual Directional Prompting, which employs multiple strategies to encourages fact-based reasoning substantially decreases the bias across all LLMs. They achieved minimal gaps, with GPT-4 exhibiting a mere 2.17% gap and 94.57% accuracy on anti sentences, demonstrate significantly greater effectiveness compared to other approaches.
Coreference Resolution with Only Syntactic Information
For Type II tasks, coreference can be resolved with only syntactic cues; thus, all models generally show higher accuracies with smaller gaps. However, since there exists a shortcut in the Type II task (the correct answer is always the second entity in the sentence), a smaller gap may not necessarily indicate less bias. Still, DDP outperforms the three baselines, particularly with GPT-4, achieving a near-negligible gap of 0.13%, suggesting that our method is highly effective in encouraging models to utilize gender-agnostic world knowledge for coreference resolution.
Summary
These experimental results suggest prompt designs that direct LLMs to rely more on gender-agnostic world knowledge (Strategy I) and less on gender shortcuts (Strategy II), thus promoting fairer and less biased decisions. Also, the performance gap between pro sentences and anti sentences decreases as the LLMs become more capable, which may indicate that LLMs are less prone to assign occupations with stereotypical gender pronouns as their general (reasoning) capabilities grow.
4.1.3 Ablation studies (WinoBias)
We further conduct ablation studies to individually examine the effectiveness of the two types of approaches, i.e., encouraging fact-based reasoning and discouraging biased reasoning. We divide the prompt Dual Directional Prompting into two parts: the Counteract Only part prompting that both occupations are equally likely to be male and female (Strategy II), and the Fact Only part informing the model which sentence(s) it regards to be more likely in real life (Strategy I).
Moreover, we categorize each LLM’s response into four groups: (i) TT: where the LLM correctly answers both the base question and the original question, (ii) TF: where the LLM only answers the base question correctly, (iii) FT: where the LLM only answers the original question correctly, and (iv) FF: where the LLM answers wrongly on both questions.
Table 2: Error Analysis on WinoBias Type I coreference task. We divide the models’ responses into 4 categories to better understand the cause of their errors and success. TT denotes the (%) of examples where LLM answers both the base question and the original question correctly, FF denotes the (%) of examples where both questions are answered incorrectly, indicating the coreference errors caused by the model’s world knowledge. TF denotes the coreference errors caused by gender bias as only base questions are correctly answered, and FT indicates coreference success that may be due to gender shortcut since the base questions get wrong but original questions are correctly answered. More details about the settings can be found in Appendix C.1.
GPT-4     DDP 94.44 96.07 1.76 0.27 0.14 0.68 3.66 2.99 Fact Only 91.45 95.93 5.29 0.14 0.54 0.81 2.71 3.12 Counteract Only 70.69 88.20 25.78 8.28 0.95 1.76 2.58 1.76 Default 69.47 87.25 26.87 8.68 1.36 2.44 2.31 1.63
As shown in Table 4.1.3, separating prompting strategies into Fact Only and Counteract Only, we observe that Fact Only still performs quite well, but Counteract Only leads to a significant decrease in performance. This suggests that while counteracting existing (selection) bias is important, providing explicit instructions that promote fact-based reasoning is crucial for mitigating bias in LLMs. Moreover, comparing the performance of GPT-4 with GPT-3.5, we can see that the GPT-4’s improvements mostly stem from better (non-gender-related) world knowledge (making way less mistakes on base questions) and self-consistency (less mistakes in the TF group). More ablation studies can be found in Appendix D.1, where we also design prompts to control the extent to which we counteract the existing bias in LLMs.
4.2 Social Bias: BBQ
The Bias Benchmark for QA (BBQ) (Parrish et al., 2021) is a dataset designed to evaluate bias for question-answering tasks. It highlight attested social biases against people belonging to protected classes along nine social dimensions relevant to U.S. English-speaking contexts. Besides common social dimensions such as age, gender, and race, BBQ also contains social categories such as physical appearance and socio-economic status where bias may be much more subtle. Each question in BBQ contains a context and 2 involved entities or persons, where the model needs to decide among 3 choices: one of the two entities, or unknown.
4.2.1 Experimental Settings (BBQ)
For our experiments, we consider the disambigous setting in BBQ where we test whether the model’s biases override a correct answer choice given an adequately informative context. In this setting, the answer is always one of the two entities rather than unknown. There are over 16,000 examples under this setting, and we measure the performance of GPT-4 across these nine social categories.